# Detecting the Root Cause Code Lines in Bug-Fixing Commits by Heterogeneous Graph Learning

### v1

@@ Abstract.
JIT Defect Prediction, SZZ, Bug-Fixing Commit
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456journal: JACMjournalvolume: 37journalnumber: 4article: 11publicationmonth: 9
@@ 1. INTRODUCTION
To improve the accuracy of this algorithm, researchers have proposed various modifications that improve B-SZZ in different dimensions. To address the noise issue in B-SZZ, Kim et al. (4, ) introduced AG-SZZ, which filters out blank lines, comment lines, and cosmetic changes in the code using an annotation graph. This enhancement allows AG-SZZ to more accurately identify true bug-inducing code commits, significantly improving the algorithm’s precision. Da Costa et al. (5, ) extended the SZZ algorithm by filtering out meta changes (such as branches, merges, and property changes), proposing MA-SZZ. Since these meta changes do not genuinely modify the source code, MA-SZZ reduces the likelihood of false positives by excluding these invalid code modifications. However, as the complexity of codebases increases, refactoring operations pose a significant challenge for the SZZ algorithm. To address this, Neto et al. (6, ) proposed RA-SZZ, which integrates refactoring detection tools, RefDiff and RefactoringMiner, to reduce false positives by identifying and excluding refactoring operations. This further optimizes RA-SZZ’s performance in complex codebases. Recently, advancements in deep learning have provided new opportunities for improving the SZZ algorithm. Neural SZZ is a deep learning-based approach (tang2023neural, ) that captures the semantic relationships between deleted lines and other code lines by constructing heterogeneous graphs of commits and using a heterogeneous graph attention network(HAN) model. This algorithm employs a learning-to-rank technique to evaluate and rank these deleted lines, offering a more intelligent and accurate approach for identifying the root causes in bug-fixing commits.
Despite the progress made by existing work, two challenges remain:
@@
Challenge 1: Struggling with Heterogeneous Commit Structures. In software engineering projects, bug-fixing commits often include many non-essential changes (21, ; 22, ), which can be highly heterogeneous and complex. Essential bug-fixing part coexists with other forms of modification, such as code refactoring, feature additions, etc. Traditional approaches either oversimplify these heterogeneous changes (assuming that all changed lines have repair characteristics: original SZZ (3, )) or separate them according to rigid predefined classification schemas for different code types or categories of changes (RA-SZZ (6, )). However, both strategies rely on static rules to solve the problem, making it difficult to pinpoint the difference between the actual fixed lines and other lines that have been changed for refactoring or enhancement.
@@
Challenge 2: Ignoring Cross-Line Dependencies in Code Changes. Semantic associations between adjacent or logically related lines are critical to understanding code changes. The functional logic of code is often scattered across multiple lines or blocks of code (23, ; shao2008understanding, ; yin2024line, ), and existing methods face challenges in capturing cross-line dependencies. Deep learning architectures often attempt to understand the ”more global” representation of lines of code through multiple layers of aggregation. However, with the increase of the range of code line interaction information, the local semantic information learned early is overshadowed by the global information, resulting in the loss of key semantic features. It is difficult to effectively capture and preserve cross-line semantic relationships, which will make the semantics of each code line tend to be homogenized  (jin2022feature, ; li2018deeper, ), and its own personality will disappear.
@@
To address these challenges, we propose a new heterogeneous graph neural network model called RC_Detector to learn the hidden semantic representations of code lines. RC_Detector consists of three main components: the bug-fixing graph construction component, the code semantic aggregation component, and the cross-line semantic retention component. Specifically, the bug-fixing graph construction component extracts the source code from both the previous and newer versions, generating the corresponding syntax trees and program dependency graphs to identify the code syntax structures and program dependencies within the bug-fixing commits. The code lines are then mapped to graph nodes based on this information, and the types of edges between nodes are determined. Finally, the source code is converted into vector representations, forming a heterogeneous graph format suitable for neural networks. The code semantic aggregation component processes each heterogeneous graph generated from bug-fixing commits by calculating the semantic similarity between the target code line node and related code line nodes, assigning weights to all related code lines accordingly. These weights are then used to aggregate the semantics of code lines directly related to the target, thereby learning the hidden semantic representation of the target code line. This component customizes the handling of different types of relationships and nodes, effectively addressing the Challenge 1. The cross-line semantic retention component calculates the attenuation and reinforcement gates using the semantic representations of the old and new code. These gates dynamically manage the flow of information by controlling the retention or updating of the original old code semantics and the aggregated new semantic representations. The RC_Detector effectively preserves cross-line semantic relationships, addressing the Challenge 2.
@@ 2. RELATED WORKS
In recent years, JIT defect prediction has transcended the limitations of traditional approaches by incorporating more sophisticated and innovative techniques. For example, Huang et al. introduced (7, ) an end-to-end deep learning framework that uses Convolutional Neural Networks (CNN) to automatically generate features from commit messages and code changes, followed by a fully connected layer for defect prediction. Building on this, Choi et al. proposed CC2Vec (8, ), which leverages a Hierarchical Attention Network (HAN) to automatically learn distributed representations of code commits, improving performance. To further enhance the efficiency of JIT defect prediction, Hoang et al. developed JITLine (9, ), an approach that combines the strengths of DeepJIT and CC2Vec, further improving prediction accuracy and granularity. Neural SZZ, proposed by Wang et al. (tang2023neural, ), extends the traditional SZZ algorithm by employing Heterogeneous Graph Neural Networks (HGNN) to capture deep semantic representations of code, focusing on identifying the root causes of defects through learning-to-rank techniques.
@@ 3.1. Overview
To accommodate the heterogeneity of code commits while accurately capturing the semantics of contextual code lines (25, ), thereby improving the defect prediction performance of JIT approaches in projects, we propose the RC_Detector model. The framework of the RC_Detector model is illustrated in Figure LABEL:fig:_framework. RC_Detector primarily consists of three components: the bug-fixing graph construction component, the code semantic aggregation component, and the cross-line semantic retention component.
@@ 3.2. Bug-Fixing Graph Construction Component
In this section, the component constructs a heterogeneous graph format suitable for neural networks by analyzing bug-fixing commits. Specifically, this process involves two sub-steps:
@@ 3.2.1. Graph Construction
For a bug-fixing commit, the first step is to extract the Java source code from the previous and newer versions. Then, use the JavaParser tool (javaparser, ) to construct their respective Abstract Syntax Trees (ASTs), referred to as ASTpre and ASTnew. The deleted lines in the bug-fixing commit are then mapped to ASTpre, marking them as deletion nodes, while the added lines are mapped to ASTnew, marking them as addition nodes. Subsequently, the static analysis tool Joern (yamaguchi2014modeling, ) is employed to construct the Program Dependency Graphs (PDGs). The dependency graph for deletion nodes is based on the previous version’s source code, and the graph for addition nodes is based on the newer version’s source code. The PDGs includes the Control Flow Graph (CFG) (allen1970control, ), Data Dependency Graph (DDG) (ferrante1987program, ) , Call Graph (CG) (ryder1979constructing, ), and Class Member Reference Graph (CMFG). Control flow or data flow edges are then added between corresponding nodes in the two versions of the PDGs using a depth-first search algorithm, generating the previous version’s graph (Gpre) and the newer version’s graph (Gnew). Finally, a line-mapping algorithm in conjunction with the GumTree tool (falleri2014fine, )is used to merge these two graphs, connecting matched deletion and addition nodes with a line mapping edge to obtain the final heterogeneous graph. The heterogeneous graph is defined as  (hu2020heterogeneous, ): ${\rm{G}}=(V,E,A,R)$, where each code line corresponds to a node $v$ satisfying $\forall v\in V$, each program dependency relation $e$ satisfying $\forall e\in E$ and the code line type mapping function $\tau(v):V\to A$ and the program dependency relation mapping function $\phi(e):E\to R$ are defined accordingly.
@@ 3.5. Pairwise Ranking of Bug-Fixing Commits
After obtaining the downstream task embedding representation of each node, we further train the model to rank bug-fixing commits. Specifically, RC_Detector utilizes the RankNet model, a pairwise ranking approach that has proven effective in real-world ranking problems (karmaker2017application, ; chapelle2011yahoo, ; song2014adapting, ). The RankNet model is trained by learning the relative priority of deleted nodes in a pairwise manner. For each bug-fixing commit, we pair it with other bug-fixing commits to obtain a series of pairs $<n_{i},n_{j}>$. RankNet processes one of the bug-fixing commit pairs at a time and obtains the corresponding scores based on the task embedding representations of the deletion nodes in each commit, denoted as $s_{i}$ and $s_{j}$. The probability that bug-fixing commit $n_{i}$ ranks higher than $n_{j}$ is then calculated accordingly:
@@ 4.1. Datasets
We conducted our experimental study using a comprehensive dataset composed of three sub-datasets containing bug-fixing and bug-inducing commits. These datasets are of higher quality and contain less noise compared to those generated by the SZZ algorithm.
@@ 4.2. State-of-The-Art Approaches
For performance comparison, we adopted the state-of-the-art approaches used by Tang et al. (tang2023neural, ), which include various machine learning algorithms: Random Forest (RF), Linear Regression (LR), Support Vector Machine (SVM), XGBoost (XGB), and K-Nearest Neighbors (KNN), as well as a deep learning-based approaches Bi-LSTM. We also employed their best state-of-the-art approach, which uses HAN as a heterogeneous graph neural network, referred to as Neural SZZ, as a benchmark for comparison. For these approaches, we attempted to reproduce their approaches on the dataset. Our goal is to determine whether RC_Detector exhibits superior performance compared to these state-of-the-art approaches. Since RC_Detector mainly compares with the existing methods with the best performance, and the original paper does not disclose the specific hyperparameter settings of these methods, we directly used the performance results reported in the original paper for all methods except Neural SZZ. For Neural SZZ, we carefully reviewed the code disclosed by the paper’s authors to confirm its completeness and consistency with the original paper description, and then performed comprehensive training on the original dataset and tried to reproduce it. Since the original author also did not give specific hyperparameters in the paper, we adjusted the parameters many times to approximate their results as closely as possible. Although the final metrics are slightly different from the original paper, the deviation is not large and may be caused by factors such as random seeds, implementation details, or version differences. For the sake of fairness and reproducibility, we still used the results reported in the original paper in subsequent comparisons.
@@ 5.2. RQ2: How does the RC_Detector impact the performance in comparison to individual components?
Motivation: This experiment focuses on two main components of the RC_Detector model: the code semantic aggregation component and the cross-line semantic retention component. The objective of this experiment is to compare the performance of models built with individual components and demonstrate that the structural framework of RC_Detector is superior to models composed of a single component.
@@
Approach: To evaluate the effectiveness of each component, we compared RC_Detector with its two variants: RC_Detector-g and RC_Detector-h. Each variant lacks a key design element. In RC_Detector-g, we removed the code semantic aggregation component and fed the input directly into the cross-line semantic retention component. In RC_Detector-h, we removed the cross-line semantic retention component and sent the output of the code semantic aggregation component directly to the RankNet model. Both RC_Detector-g and RC_Detector-h share the same graph construction process as RC_Detector.
@@
Results: Table 4 compares the performance of RC_Detector with its two variants, RC_Detector-g and RC_Detector-h, in identifying bug-inducing commits. The best results are highlighted in bold. Except for Recall@3 and MFR, RC_Detector outperforms both variants across all metrics. In Recall@1, RC_Detector surpasses RC_Detector-g and RC_Detector-h by 1.7% and 4.9%, respectively. Similarly, in Recall@2, it exceeds the variants by 2.5% and 3.8%. Although RC_Detector does not achieve the best results in Recall@3 and MFR, the differences between its performance and the best results from RC_Detector-g are minimal, with a decrease of only 0.6% and 6%, respectively, while it outperforms RC_Detector-h by 3.1% and 16.3%. The performance differences can be attributed to the design of the model components. Unlike RC_Detector, RC_Detector-g only considers the semantic representation of individual code lines and applies gating to regulate this representation, neglecting the influence of context, which results in inferior performance on several metrics. For RC_Detector-h, although it accounts for the influence of related code lines, its lack of regulating capability for propagated information prevents it from effectively retaining cross-line semantic relationships and critical information. That leads to its weaker performance compared to RC_Detector. Additionally, we observed that the performance of RC_Detector-g consistently surpasses that of RC_Detector-h, which aligns with the findings of Tang et al. (tang2023neural, ). The semantic representation of individual code lines plays a more crucial role in identifying bug-inducing commits than that of related code lines. The advantage of RC_Detector over models composed of a single component lies in its ability to combine multiple components synergistically, thereby better capturing the contextual semantics of code lines and improving the accuracy of JIT prediction. However, compared to single-component models, RC_Detector’s drawback is that it requires more computational resources and time.
@@
Conclusion: The experimental results show that the RC_Detector model outperforms models composed of single components in terms of Recall@N and MFR. Thus, the structural framework of RC_Detector offers greater potential for improvement than individual components. This makes RC_Detector an effective approach for identifying the root causes of bugs, thereby enhancing JIT defect prediction.

### v2

@@ Abstract.
Heterogeneous Graph Learning, JIT Defect Prediction, Bug-Fixing Commit
copyright: acmcopyrightjournalyear: 2025doi: 10.1145/1122445.1122456journal: TOITjournalvolume: 1journalnumber: 2article: 3publicationmonth: 5
@@ 1. INTRODUCTION
To improve the accuracy of this algorithm, researchers have proposed various modifications that improve B-SZZ in different dimensions. To address the noise issue in B-SZZ, Kim et al. (4, ) introduced AG-SZZ, which filters out blank lines, comment lines, and cosmetic changes in the code using an annotation graph. This enhancement allows AG-SZZ to more accurately identify true bug-inducing code commits, significantly improving the algorithm’s precision. Da Costa et al. (5, ) extended the SZZ algorithm by filtering out meta changes (such as branches, merges, and property changes), proposing MA-SZZ. Since these meta changes do not genuinely modify the source code, MA-SZZ reduces the likelihood of false positives by excluding these invalid code modifications. However, as the complexity of codebases increases, refactoring operations pose a significant challenge for the SZZ algorithm. To address this, Neto et al. (6, ) proposed RA-SZZ, which integrates refactoring detection tools, RefDiff and RefactoringMiner, to reduce false positives by identifying and excluding refactoring operations. This further optimizes RA-SZZ’s performance in complex codebases.
@@
Despite the progress made by existing work, one challenge remains: In software engineering projects, bug-fixing commits often include many non-essential changes (21, ; 22, ), which can be highly heterogeneous and complex. Essential bug-fixing part coexists with other forms of modification, such as code refactoring, feature additions, etc. Traditional approaches either oversimplify these heterogeneous changes (assuming that all changed lines have repair characteristics: original SZZ (3, )) or separate them according to rigid predefined classification schemas for different code types or categories of changes (RA-SZZ (6, )). However, both strategies rely on static rules to solve the problem, making it difficult to pinpoint the difference between the actual fixed lines and other lines that have been changed for refactoring or enhancement.
Neural SZZ was proposed by Tang et al. (tang2023neural, )to solve this problem, Neural SZZ is a deep learning-based approach that captures the semantic relationships between deleted lines and other logically related lines by constructing heterogeneous graphs of commits and using a heterogeneous graph attention network(HAN) model. This algorithm evaluates and sorts based on the semantics of these deleted lines, offering a more intelligent and accurate approach for identifying the root causes in bug-fixing commits.
@@
Using Semantic associations between adjacent or logically related lines are critical to understanding code changes, the functional logic of code is often scattered across multiple lines or blocks of code (23, ; shao2008understanding, ; yin2024line, ). Deep learning architectures often attempt to understand the ”more global” representation of lines of code through multiple layers of aggregation. However, with the increase of the range of code line interaction information, the local semantic information learned early is overshadowed by the global information, resulting in the loss of key semantic features. It is difficult to effectively capture and preserve cross-line semantic relationships, which will make the semantics of each code line tend to be homogenized  (jin2022feature, ; li2018deeper, ), and its own personality will disappear.
@@
To address the challenge, we propose a new heterogeneous graph neural network model called RC_Detector to learn the hidden semantic representations of code lines. RC_Detector consists of three main components: the bug-fixing graph construction component, the code semantic aggregation component, and the cross-line semantic retention component. Specifically, the bug-fixing graph construction component extracts the source code from both the previous and newer versions, generating the corresponding syntax trees and program dependency graphs to identify the code syntax structures and program dependencies within the bug-fixing commits. The code lines are then mapped to graph nodes based on this information, and the types of edges between nodes are determined. Finally, the source code is converted into vector representations, forming a heterogeneous graph format suitable for neural networks. The code semantic aggregation component processes each heterogeneous graph generated from bug-fixing commits by calculating the semantic similarity between the target code line node and related code line nodes, assigning weights to all related code lines accordingly. These weights are then used to aggregate the semantics of code lines directly related to the target, thereby learning the hidden semantic representation of the target code line. The cross-line semantic retention component calculates the attenuation and reinforcement gates using the semantic representations of the old and new code. These gates dynamically manage the flow of information by controlling the retention or updating of the original old code semantics and the aggregated new semantic representations to preserve cross-line semantic relationships.
The bug-fixing graph construction and code semantic aggregation components draw on the existing heterogeneous graph neural network technology  (tang2023neural, ), and we improve the code semantic aggregation components. Our main contribution is the introduction of the cross-line semantic retention component, which addresses the deficiency of the existing model in preserving cross-line semantic relations during bug fixes.
@@ 2. RELATED WORKS
In recent years, JIT defect prediction has transcended the limitations of traditional approaches by incorporating more sophisticated and innovative techniques. For example, Huang et al. introduced (7, ) an end-to-end deep learning framework that uses Convolutional Neural Networks (CNN) to automatically generate features from commit messages and code changes, followed by a fully connected layer for defect prediction. Building on this, Choi et al. proposed CC2Vec (8, ), which leverages a Hierarchical Attention Network (HAN) to automatically learn distributed representations of code commits, improving performance. To further enhance the efficiency of JIT defect prediction, Hoang et al. developed JITLine (9, ), an approach that combines the strengths of DeepJIT and CC2Vec, further improving prediction accuracy and granularity. Neural SZZ, proposed by Tang et al. (tang2023neural, ), extends the traditional SZZ algorithm by employing Heterogeneous Graph Neural Networks (HGNN) to capture deep semantic representations of code, focusing on identifying the root causes of defects through learning-to-rank techniques. Neural SZZ does not rely solely on superficial textual differences or syntactic rules, but adds processing related to semantic features. First, syntactic features are obtained using an abstract syntax tree. Second, a vectorised representation of the source code is obtained, and a deep learning model is used to capture semantic features. Finally, a ranking algorithm is used to recommend the line that is the root cause of the bug.
@@ 3.1. Overview
To accommodate the heterogeneity of code commits while accurately capturing the semantics of contextual code lines (25, ), thereby improving the defect prediction performance of JIT approaches in projects, we propose the RC_Detector model. The framework of the RC_Detector model is illustrated in Fig 1. RC_Detector primarily consists of three components: the bug-fixing graph construction component, the code semantic aggregation component, and the cross-line semantic retention component.
@@ 3.2. Bug-Fixing Graph Construction Component
In this section, the component constructs a heterogeneous graph format suitable for neural networks by analyzing bug-fixing commits (tang2023neural, ). Specifically, this process involves two sub-steps:
@@ 3.2.1. Graph Construction
For a bug-fixing commit, the first step is to extract the Java source code from the previous and newer versions. Then, use the JavaParser tool (javaparser, ) to construct their respective Abstract Syntax Trees (ASTs), referred to as ASTpre and ASTnew. Subsequently, the static analysis tool Joern (yamaguchi2014modeling, ) is employed to construct the Program Dependency Graphs (PDGs). The PDGs includes the Control Flow Graph (CFG) (allen1970control, ), Data Dependency Graph (DDG) (ferrante1987program, ) , Call Graph (CG) (ryder1979constructing, ), and Class Member Reference Graph (CMFG). Control flow or data flow edges are then added between corresponding nodes in the two versions of the PDGs using a depth-first search algorithm, generating the previous version’s graph (Gpre) and the newer version’s graph (Gnew). Finally, a line-mapping algorithm in conjunction with the GumTree tool (falleri2014fine, )is used to merge these two graphs, connecting matched deletion and addition nodes with a line mapping edge to obtain the final heterogeneous graph. The heterogeneous graph is defined as  (hu2020heterogeneous, ): ${\rm{G}}=(V,E,A,R)$, where each code line corresponds to a node $v$ satisfying $\forall v\in V$, each program dependency relation $e$ satisfying $\forall e\in E$ and the code line type mapping function $\tau(v):V\to A$ and the program dependency relation mapping function $\phi(e):E\to R$ are defined accordingly.
@@ 3.5. Pairwise Ranking of Bug-Fixing Commits
After obtaining the downstream task embedding representation of each node, we further train the model to rank bug-fixing commits. Specifically, RC_Detector utilizes the RankNet model same to NeuralSZZ  (tang2023neural, ). RankNet is a pairwise ranking approach that has proven effective in real-world ranking problems (karmaker2017application, ; chapelle2011yahoo, ; song2014adapting, ). The RankNet model is trained by learning the relative priority of deleted nodes in a pairwise manner. For each bug-fixing commit, we pair it with other bug-fixing commits to obtain a series of pairs $<n_{i},n_{j}>$. RankNet processes one of the bug-fixing commit pairs at a time and obtains the corresponding scores based on the task embedding representations of the deletion nodes in each commit, denoted as $s_{i}$ and $s_{j}$. The probability that bug-fixing commit $n_{i}$ ranks higher than $n_{j}$ is then calculated accordingly:
@@ 4.1. Datasets
We conducted our experimental study using a comprehensive dataset  (tang2023neural, )composed of three sub-datasets containing bug-fixing and bug-inducing commits. These datasets are of higher quality and contain less noise compared to those generated by the SZZ algorithm.
@@ 4.2. State-of-The-Art Approaches
For performance comparison, we adopted the state-of-the-art approaches used by Tang et al. (tang2023neural, ), which include various machine learning algorithms: Random Forest (RF), Linear Regression (LR), Support Vector Machine (SVM), XGBoost (XGB), and K-Nearest Neighbors (KNN), as well as a deep learning-based approach Bi-LSTM. We also employed their state-of-the-art approach, which uses HAN as a heterogeneous graph neural network, referred to as Neural SZZ, as a baseline for comparison. For these approaches, we attempted to reproduce their approaches on the dataset. Our goal is to determine whether RC_Detector exhibits superior performance compared to these state-of-the-art approaches. Since RC_Detector mainly compares with the existing methods with the best performance, and the original paper does not disclose the specific hyperparameter settings of these methods, we directly used the performance results reported in the original paper for all methods except Neural SZZ. For Neural SZZ, we carefully reviewed the code disclosed by the paper’s authors to confirm its completeness and consistency with the original paper description, and then performed comprehensive training on the original dataset and tried to reproduce it. Since the original author also did not give specific hyperparameters in the paper, we adjusted the parameters many times to approximate their results as closely as possible. Although the final metrics are slightly different from the original paper, the deviation is not large and may be caused by factors such as random seeds, implementation details, or version differences. For the sake of fairness and reproducibility, we still used the results reported in the original paper in subsequent comparisons.
@@ 5.2. RQ2: How does the RC_Detector impact the performance in comparison to individual components?
Motivation: This experiment focuses on two main components of the RC_Detector model: the code semantic aggregation component and the cross-line semantic retention component. The objective of this experiment is to compare the performance of models built with individual components and demonstrate that the structural framework of RC_Detector is superior to models composed of a single component. Meanwhile, to further examine the role of the cross-line semantic retention component, we conducted the experiment of Neural SZZ-GRU, which is the original Neural SZZ method plus the cross-line semantic retention component.
@@
Approach: To evaluate the effectiveness of each component, we compared RC_Detector with its two variants: RC_Detector-g and RC_Detector-h. Each variant lacks a key design element. In RC_Detector-g, we removed the code semantic aggregation component and fed the input directly into the cross-line semantic retention component. In RC_Detector-h, we removed the cross-line semantic retention component and sent the output of the code semantic aggregation component directly to the RankNet model. Both RC_Detector-g and RC_Detector-h share the same graph construction process as RC_Detector. We also conducted an experiment to compare the situations where the Neural SZZ method used GRU and did not use GRU.
@@
Results: Table 4 compares the performance of RC_Detector with its two variants, RC_Detector-g and RC_Detector-h, in identifying bug-inducing commits. The best results are highlighted in bold. Except for Recall@3 and MFR, RC_Detector outperforms both variants across all metrics. In Recall@1, RC_Detector surpasses RC_Detector-g and RC_Detector-h by 1.7% and 4.9%, respectively. Similarly, in Recall@2, it exceeds the variants by 2.5% and 3.8%. Although RC_Detector does not achieve the best results in Recall@3 and MFR, the differences between its performance and the best results from RC_Detector-g are minimal, with a decrease of only 0.6% and 6%, respectively, while it outperforms RC_Detector-h by 3.1% and 16.3%. The performance differences can be attributed to the design of the model components. Unlike RC_Detector, RC_Detector-g only considers the semantic representation of individual code lines and applies gating to regulate this representation, neglecting the influence of context, which results in inferior performance on several metrics. For RC_Detector-h, although it accounts for the influence of related code lines, its lack of regulating capability for propagated information prevents it from effectively retaining cross-line semantic relationships and critical information. That leads to its weaker performance compared to RC_Detector. Additionally, we observed that the performance of RC_Detector-g consistently surpasses that of RC_Detector-h, which aligns with the findings of Tang et al. (tang2023neural, ). The semantic representation of individual code lines plays a more crucial role in identifying bug-inducing commits than that of related code lines. The advantage of RC_Detector over models composed of a single component lies in its ability to combine multiple components synergistically, thereby better capturing the contextual semantics of code lines and improving the accuracy of JIT prediction. However, compared to single-component models, RC_Detector’s drawback is that it requires more computational resources and time. Meanwhile, the cross-line semantic retention component has also achieved significant improvements in the Neural SZZ method. Using GRU has increased by 0.6%, 3.4%, 2.8%, and 23. 8%, respectively, in various indicators compared to not using it.
@@
Conclusion: The experimental results show that the RC_Detector model outperforms models composed of single components in terms of Recall@N and MFR. Thus, the structural framework of RC_Detector offers greater potential for improvement than individual components. This makes RC_Detector an effective approach for identifying the root causes of bugs, thereby enhancing JIT defect prediction. The results also show that the use of the GRU part is always much better than when it is not used, proving the importance of the cross-line semantic retention component.